Solve each system by the addition method.
(1.5)x-2y=-0.25
3x+(1.5)y=6.375

Answer by checkley71(8403)

MULTIPLY THE FIRST EQUATION BY -2 & ADD THEM
-3X+4Y=.5
3X+1.5Y=6.375
--------------------------
5.5Y=6.875
Y=6.875/5.5
Y=1.25 ANSWER
3X+1.5*1.25=6.375
3X+1.875=6.375
3X=6.375-1.875
3X=4.5
X=4.5/3
X=1.5 ANSWER.
PROOF:
1.5*1.5-2*1.25=-.25
2.25-2.5=-.25
-.25=-.25